OpenStudy (anonymous):

Which function does NOT have an inverse? A)f(x) = 2x + 10, x ≥ -5 B)f(x) = 2x + 10, x ≥ 0 C)f(x) = 2x2 + 10, x ≥ -5 D)f(x) = 2x2 + 10, x ≥ 0

OpenStudy (anonymous):

Okay, first things first, linear functions always have inverses, so scratch A and B.

OpenStudy (anonymous):

okay the last two are squared btw they didnt show up

OpenStudy (anonymous):

Yeah... they both have the same function, but different domains... that's gotta be tough ^.^

OpenStudy (anonymous):

i needed an explination

OpenStudy (anonymous):

not sarcasm.

OpenStudy (anonymous):

f(x) = 2x^2 + 10 \[\large y = 2x^2 + 10\]So what I want you to do is switch x and y \[\large x = 2y^2 + 10\] And make y alone.

OpenStudy (anonymous):

isnt it \[\sqrt{5-x}=y\]

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

first i subtracted 10

OpenStudy (anonymous):

10-x=2y\[10-x=2y^{2}\]

OpenStudy (anonymous):

2nd one

OpenStudy (anonymous):

then i divided both sides by 2

OpenStudy (anonymous):

\[10-x \div 2 =y ^{2}\]

OpenStudy (anonymous):

then i square rooted both sides

OpenStudy (anonymous):

for a function to have an inverse it must be one-one and onto type. I trust you know the two requirements?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, put your set of requirements, the correct out to be D because of onto type of function.

OpenStudy (anonymous):

I mean, D ought to be invertible. *pardon*

OpenStudy (anonymous):

@PeterPan Inverting a function \(f(x)\), you find \(f^{-1}(x)\)

OpenStudy (anonymous):

C is not of onto type, and thus it is a non-invertible function. (I may be wrong here, relations and functions is not one of my better topics)

OpenStudy (anonymous):

@natalia517 you are on the right track. follow @NeetziD

OpenStudy (anonymous):

yeah, and that still involves switching x and y, and then isolating y. The resulting function of x would be f^-1(x)

OpenStudy (anonymous):

alright thank you for helping me again neetziD

OpenStudy (anonymous):

Don't mention. ;)

OpenStudy (anonymous):

check: for option C) \[f(x)=y=2x^2+10,\qquad\qquad\forall\quad x\ge-5\text{and}\;0\le y<\infty\\ \frac{y-10}{2}=x^2\implies x=\pm\sqrt{y-10\over2} \] this limits the possible range for "y"...

OpenStudy (anonymous):

but you will notice that the domain of \(f(x)\) is \(10\le y<\infty\)

OpenStudy (anonymous):

i meant, the "range"

OpenStudy (anonymous):

and for the inverse, it is no longer one-to-one correcpondence and the difinition of function gets mangled. so cannot be called a function anymore for the D) option, though it looks just the same, "x"can take only the POSITIVE values and maintains the one-to-one correspondence. and its inverse is a valid function.

OpenStudy (anonymous):